OpenStudy (anonymous):

MEDAL!!!!!! Solve for n. –9 = 5/17 n

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

–9 = 5/17 n -9*17n=5

OpenStudy (anonymous):

-153n=5 n =-5/153

OpenStudy (anonymous):

thanks so much! can you solve this one too please? 3 + 6z = 13 + 6z Solve for z.

OpenStudy (anonymous):

3-13=6z-6z -10=0 u must check the statement there must be + , - mistake

OpenStudy (anonymous):

no, that's the right problem. here are the options for the answers: 3 + 6z = 13 + 6z z = – 5/6 z = 2 2/3 z = infinitely many solutions z = no solution

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

thanks so much, this is the last problem i don't get. do you think you could answer it? Which equation has no solution? 8 – (5v + 3) = 5v – 5 3m – 6 = 5m + 7 – m 3w + 4 – w = 5w – 2(w – 2) 7y + 9 = 7y – 6 if you do i'll become your fan as well as give you a medal :) its okay if you can't though

OpenStudy (anonymous):

simplify each equation, and the equation with no solution is the one that does not make sense when it is simplified an example would be "0=1"

OpenStudy (anonymous):

i know what to do to answer it, i just don't know how to simply those :/

OpenStudy (anonymous):

*simplify

OpenStudy (anonymous):

when we are not able to find any value for x, v or any variable, then it means it has no solution

OpenStudy (anonymous):

okay then, one sec. i'll try...

OpenStudy (anonymous):

i think it's the second one... am i right?

OpenStudy (anonymous):

umm anyways i'll jus ask this in a seperate question. thanks guys!

OpenStudy (anonymous):

its not the second one please show steps